Given below is the TC of a firm
TC=100+50q -12q^2+q^3
Calculate the level of output
1
Expert's answer
2020-12-29T15:39:54-0500
"TR=pq"

"MR=p"


"MC=50-24q+3q^2"

"MC=MR"

"3q^2-24q+50-p=0"

"q=4+\frac{\sqrt{3p-6}}{3} or q=4-\frac{\sqrt{3p-6}}{3}"